f(x)=4x-3 and g(x)=(x+3)/4
a) Find the function fg(x)
Hence describe the relationship between the functions f and g.
b) Write down the exact value of fg(sqrt(3))

Answer by stanbon(75887)

f(x)=4x-3 and g(x)=(x+3)/4
a) Find the function fg(x)
----
Comment: I'm going to assume you mean fog(x)
------------------------------------------------
fog(x) = f[(x+3)/4] = 4[(x+3)/4] - 3 = x+3-3 = x
-----------------------------------------------------
Hence describe the relationship between the functions f and g.
Ans: f and g are inverse functions---each function undoes what
the other function does.
--------------------------------------
b) Write down the exact value of fog(sqrt(3)) = sqrt(3)
You feed sqrt(3) to g and it gives you g(sqrt(3))
You then feed g(sqrt(3)) to f and it gives you back sqrt(3)
--------------
Cheers,
Stan H.